OpenStudy (anonymous):

ok im havin a issue here i get 760 miles on 30.4 gallons of gas whats my miles a gallon im thinking 25

OpenStudy (anonymous):

Miles per gallon would be 760/30.4 = 25

OpenStudy (anonymous):

so i was rite with 25 miles per gallon?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i new it thanks for the help can youdo one more for me i just wanna check my answers

OpenStudy (anonymous):

If we get 340 miles on 16 gallons of gas, what is the miles per gallon?

OpenStudy (anonymous):

Try that one.

OpenStudy (anonymous):

Hint: The answer is in units (miles per gallons), which is \[miles/galon\]

OpenStudy (anonymous):

21.4/a gallon?

OpenStudy (anonymous):

The answer is (340 miles) / (16 gallons) = 85/4 = 21.25

OpenStudy (anonymous):

Your answer is alright. Good job.

OpenStudy (anonymous):

great i have trouble sometimes

OpenStudy (anonymous):

Do you understand this?

OpenStudy (anonymous):

yes that one i did

OpenStudy (anonymous):

may I ask you another and give you the answer i got

OpenStudy (anonymous):

A car travels 625 miles on 25 miles of gas. What is the miles per gallon for that car?

OpenStudy (anonymous):

26

OpenStudy (anonymous):

The answer is (miles traveled) / (gallons of gas) because its units are miles/gallon. You are very close with 26.

OpenStudy (anonymous):

Try again.

OpenStudy (anonymous):

28/1 gallon

OpenStudy (anonymous):

I got 25, because 625/25 = 25